OpenStudy (anonymous):

Write the equation of the line that passes through the points (12, 6) and (12,-4).

OpenStudy (anonymous):

Use two point formula: \[\large \frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x _2 - x_1}\]